OpenStudy (anonymous):

Tracy purchased a building for $209,350. The land appreciates about 4.5% each year. What is the value of the land after 15 years?

OpenStudy (jlg030597):

ok what do you do first i know what to do im asking you.

OpenStudy (anonymous):

I have no idea.

OpenStudy (jlg030597):

ok let me do the problem really quick so i have the answer

OpenStudy (jlg030597):

@ikram002p I need some help here its been a while since I did this

OpenStudy (jlg030597):

lets get some help

OpenStudy (jlg030597):

We need some help lol

OpenStudy (anonymous):

yes we do lol

OpenStudy (mathmale):

this problem is done in the same way as is the calculation of the Amount of money one would have after investing Principal P at interest rate r for n years. What is P in this case? r? t?

OpenStudy (mathmale):

what does the pertinent equation / formula look like?

OpenStudy (anonymous):

A = (Prt)+P

OpenStudy (ikram002p):

the binifit each year = $209,350 ×4.5% so since its 15 yrs it would be = $209,350 ×4.5% ×15 the value of the land after 15 years =binefit+the orginal

OpenStudy (jlg030597):

so I got it right crap i second guessed my self I know the answer now

OpenStudy (ikram002p):

no @jlg030597 u did ur best to help :D

OpenStudy (mathmale):

or, A = P (1+r)^t

OpenStudy (jlg030597):

no I mean I did that with the multiplying everything and got the answer from that or no?

OpenStudy (jlg030597):

and thanks @ikram002p

OpenStudy (ikram002p):

np ur wlc anytime :D

OpenStudy (jlg030597):

I think I got it right Atleast but im not sure would the answer be in the millions

OpenStudy (ikram002p):

it would be for sure

OpenStudy (anonymous):

The answer i got isnt one of them opition

OpenStudy (jlg030597):

ok I may have gotten it then

OpenStudy (jlg030597):

what did you get?

OpenStudy (anonymous):

141 311.25

OpenStudy (jlg030597):

are you sure its not 14131125 not with a decimal

OpenStudy (anonymous):

$221,440.14 $104,935.77 $405,151.38 $350,661.25 these are my opition

OpenStudy (jlg030597):

ok wait a second

OpenStudy (mathmale):

Why not write out the calculation that you need to perform? A = value of the land after 15 years = P*(1+r)^t A = $209350 (1 + 0.045)^15 Please re-calculate this. If you do, and if you do it correctly, you'll immediately have the correct answer.

OpenStudy (mathmale):

what kind of calculator have you?

OpenStudy (anonymous):

im using google m calculator ran out of batteries

OpenStudy (anonymous):

405 151.379

OpenStudy (mathmale):

OK. Type in 209350(1+0.045)^15 and see what Google gives you.

OpenStudy (anonymous):

405 151.379